OpenStudy (lgbasallote):

radium disintegrates at a rate proportional to the amount present. if 100 mg are set aside now, 96 mg will be left 100 years later. find how much will be left 2.56 centuries later and the half life too.

OpenStudy (lgbasallote):

do i use \[100 - 4^t?\] \[\implies 100 - 4^{2.56} = 65.22 = 66\] lol what did i do wrong @satellite73 ?

OpenStudy (anonymous):

\[\frac{96}{100}=.96\] so model as \[100\times(.96)^{\frac{t}{100}}\]

OpenStudy (anonymous):

you need the proportion it decreases by, not the absolute number decreases by 4% retains 96% of its current amount

OpenStudy (lgbasallote):

this is pretty tricky haha

OpenStudy (anonymous):

half live is easy enough solve \[.96^{\frac{t}{100}}=\frac{1}{2}\] for \(t\) of course you will need logs for this solution

OpenStudy (lgbasallote):

hmm you dont use 0,5 instead of 0,96?

OpenStudy (anonymous):

but it is a on step solution \[t=100\times \frac{\log(.5)}{\log(.96)}\]

OpenStudy (lgbasallote):

it's lacking one year :O t = 1697.9 it should be 1699

OpenStudy (anonymous):

ok i have confused you let me back up a step you want the half life. you start with 100 so when half is gone you will have 50 you want to solve \[100\times (.96)^{\frac{t}{100}}=50\] for \(t\) but the first step is to divide both sides by 100 and write \[(.96)^{\frac{t}{100}}=\frac{1}{2}\]

OpenStudy (lgbasallote):

oh i see what happened

OpenStudy (anonymous):

i get 1698 rounded

OpenStudy (anonymous):

of course it doesn't matter what you start with, half is still half so you always set it equal to \(\frac{1}{2}\)

OpenStudy (anonymous):

i am not saying you shouldn't learn how to do it the other way, i am just saying this method is intuitive, snappy, and more accurate because if you solve for \(k\) and use \(p_0e^{kt}\) then since \(k\) will be a log of some sort, you will no doubt truncate the decimal when you solve and therefore be off by a little

OpenStudy (lgbasallote):

hmm i tried using diff eq but it gave that too

OpenStudy (anonymous):

other method you need to solve \[e^{100k}=.96\] for \(k\) not hard, you get \[k=\frac{\log(.96)}{100}=-0.0004082199452025517\] then write the equation as \(100e^{-0.0004082199452025517t}\)

OpenStudy (lgbasallote):

how come when i do ln2/0.0408 then it gives the right answer?

OpenStudy (lgbasallote):

the more i add decimal places the lower the value

OpenStudy (anonymous):

i do not know where you got those numbers from

OpenStudy (lgbasallote):

\[\ln (\frac{100}{96}) = 0.0408\]

OpenStudy (anonymous):

you have it upside down

OpenStudy (anonymous):

\[100\times e^{100t}=96\] \[e^{100t}=.96\] \[100t=\log(.96)\] \[t=\frac{\log(.96)}{100}\]

OpenStudy (anonymous):

and the reason you got the right answer is that you had the numerator upside down as well should be \(\log(\frac{1}{2})\)

OpenStudy (anonymous):

are you using centuries for your unit, or years? looks like you might be using centuries, which is fine in that case you solve \[e^k=.96\] and get \[k=\log(.96)=-.0408\]

OpenStudy (anonymous):

then to solve for the half life put \[e^{-0.408t}=\frac{1}{2}\] \[t=\frac{\ln(\frac{1}{2})}{-.0408}\]

OpenStudy (anonymous):

now your answer is in centuries instead of years

OpenStudy (anonymous):

oops typo there, first line should have been \[e^{-0.0408t}=\frac{1}{2}\]

OpenStudy (lgbasallote):

but if i use e^-0.0408 = 1/2 then wont it be 1.92?

OpenStudy (anonymous):

not according to wolfram http://www.wolframalpha.com/input/?i=e^%28-.0408t%29%3D.5

OpenStudy (anonymous):

should be \(t=\frac{\log(.5)}{-.0408}\)